OpenStudy (anonymous):

Find the lowest \[N \in \mathbb{N} \] meeting the requirements: \[|\frac{ n-7 }{ n+7 }-1|\le \frac{ 1 }{ 1000 } \forall n \ge N\]

ganeshie8 (ganeshie8):

\(\left|\dfrac{n-7}{n+7}-1\right|\) \(=\left|\dfrac{(n+7)-14}{n+7}-1\right|\) \(=\dfrac{14}{n+7}\) You want this expression to not exceed \(\frac{1}{1000}\) : \(\dfrac{14}{n+7}\le \dfrac{1}{1000}\) see if you can solve \(n\)

OpenStudy (anonymous):

If i solve n, i get: \[n \ge 13993, n <-7\] What does this tell me of N?

ganeshie8 (ganeshie8):

you're given that \(N\in\mathbb{N}\) so we only consider positive values of \(n\)

OpenStudy (anonymous):

\[\frac{ 1 }{ \frac{ n+7 }{ 14 } }\le \frac{ 1 }{ 1000 }\] \[\frac{ n+7 }{ 14 }\ge1000\] \[n+7\ge14000,n \ge 13993\]

OpenStudy (anonymous):

Ohh yes, that is correct. That gives the interval: \[[13993, infinity)\]

ganeshie8 (ganeshie8):

\(n \ge 13993\) tells us that the given statement is true when ever \(n\) is at least \(13993\). In other words \(N = 13993\)

OpenStudy (anonymous):

But I need to find the lowest N, and if \[n \ge N\], couldnt N be 1 then? In this case it would still be lower than n and in the natural numbers.

ganeshie8 (ganeshie8):

\(N\) is the lowest \(n\) that satisfies the given inequality.

ganeshie8 (ganeshie8):

since \(13993\) is the lowest \(n\) that satisfies the given inequality, we have \(N = 13993\)

OpenStudy (anonymous):

Ohh, okay. That makes sense, thank you so much :)

ganeshie8 (ganeshie8):

np :)